• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 00:59:29 on March26,2014

    Entry number 387566

    keyword=BPM pedestal calibration

    We did some pulser runs with no beam to check BPM pedestals,
    since Luke fixed the noise problem on the L-HRS today. He fixed it
    by relocating the offset/attenuator boards. The speculation is that
    there is a switching power supply on the 2nd level which caused the noise.

    Runs 1241 and 20255 are with forced gain=0 and T8 trigger

    Runs 1242 and 20256 are with autogain and T8 trigger

    Runs 1243 and 20257 are returned to cosmics.

    The pedestals are all around 3500 with RMS of about 10. Typical ADC spectra
    in fig 1 and 2.

    Pedestals for channels 1-8 (in order they are plugged in) are as follows. I'll put
    these into the BPM database

    L-HRS
    3708, 3534, 3483, 3475, 3456, 3393, 3401, 3278

    R-HRS
    3510, 3505, 3579, 3662, 3563, 3462, 3330, 3546

    The raster pick-off also has pedestals of around this size and width but
    I don't think these get used by the software.


    FIGURE 1

    FIGURE 2